# Forgetting-MarI: LLM Unlearning via Marginal Information Regularization

### v2

@@ 1 Introduction
As machine learning models, particularly Large Language Models (LLMs), get trained on bigger datasets containing potentially sensitive or regulated information, and as LLMs are increasingly deployed in high-stakes domains, the need to selectively remove specific data influences from these models has become critical. This requirement is driven not only by privacy regulations such as the European Union’s General Data Protection Regulation (GDPR) and its “right to be forgotten,” but also by practical concerns including the removal of copyrighted content, personally identifiable information, or data determined to be harmful or biased [metz2024openai, grynbaum2023times, freeman2024exploring, cyphert2023generative, carlini2021extract]. Unlearning, or removing the influence of specific data post hoc, is an attractive tool for achieving this information removal, especially with the high costs of retraining a model from scratch.
@@ 1.1 Open Challenges in LLM Unlearning
Effective LLM unlearning must balance three objectives [liu2024rethinkingmachineunlearninglarge]. First, unlearn efficacy measures how well a model suppresses the influence of the data we want to unlearn, called the unlearn set $\mathcal{D}_{u}$. Second, utility preservation ensures the model’s ability to retain performance on general tasks and the data we are still authorized to use, called the retain set, $\mathcal{D}_{r}$ is not lost. Finally, computational cost encompasses the time, memory, and carbon used during unlearning. All unlearning techniques aim to optimize these three objectives, which inherently come with tradeoffs; what differs is where and how the model parameters are updated, directly affecting their ability to balance the three. A breakdown of existing techniques and their strengths and weaknesses is shown in Table 1, with their technical details and commonality in indirect marginal unlearning in Appendix A.1.
@@
Numbers map to BibTeX entries: 1[yao2024llmunlearning], 2[liu2024rethinkingmachineunlearninglarge], 3[hong2024dissect], 4[rafailov2023dpo], 5[zhang2024npo], 7[meng2022rome], 8[meng2023memit], 9[fang2025alphaedit], 10[wu2023depn], 11[ishibashi2024sanitization], 12[eldan2023approximate], 13[shi2024ulmr], 14[liu2024sku], 15[ilharco2023task], 16[gao2024o3].
@@
Robust unlearning & Utility Preservation: Existing LLM unlearning techniques via full-parameter fine-tuning typically treat the unlearn set $\mathcal{D}_{u}$ as fully toxic, forcing the model to forget every sequence in $\mathcal{D}_{u}$ regardless of their overlap with the retain set $\mathcal{D}_{r}$. Examples include loss-reversal [liu2022continual], gradient-difference [yao2024llmunlearning], KL-ascent [hong2024dissect], and preference-based DPO/NPO [rafailov2023dpo, zhang2024npo]. Even local editors that aim to make precise edits (ROME, MEMIT) share this limitation [meng2022rome, meng2023memit], erasing shared facts and stylistic cues, and raising perplexity on $\mathcal{D}_{r}$ and held-out tasks. Benchmarks (RWKU, MUSE, Eight-Method) consistently report sizable utility drops after unlearning [jin2024rwku, muse2024benchmark, lynch2024eight].
Stable Continual Unlearning: As the legal landscape around data usage changes, a deployed LLM may receive hundreds or thousands of unlearn requests. Production-ready unlearning, therefore, needs to be able to repeatedly unlearn, retain utility, and keep computation and memory within a practical range. Exact methods like full retraining or shared SISA guarantee unlearning but their cost scales with both model size and request count [cao2015machine, ginart2019unlearning, bourtoule2021sisa]. Lighter updates like influence functions [guo2020certified] or repeated ROME/MEMIT edits [meng2022rome, meng2023memit] are cheap per removal yet accumulate inference costs and utility drift. Task-vector subtraction or adapter stacks save compute during unlearning but require storing external model adapters, also creating downstream inference costs [ilharco2023task, gao2024o3, wang2024ood]. Thus, continually unlearning without runaway resources or utility loss remains unsolved.
Formal Guarantees at LLM-Scale: Certified unlearning is well established for linear/kernel models [guo2020certified], high-dimensional classifiers [ding2024highdim], and general mathematical formulations of machine unlearning [xu2025machine]. However, no existing method provides guarantees that scale to autoregressive transformers with billions of parameters (7B–70B+), such as GPT or Llama. As a result, practitioners lack reliable guarantees of the extent to which the unlearn set remains uninferable or undetectable after common downstream operations such as compression, distillation, or adversarial probing [liu2024rethinkingmachineunlearninglarge].
@@ 2.2 Marginal Information & Perplexity-Based Detectors
be the standard cross-entropy (per-token negative log-likelihood). State-of-the-art detectors [carlini2021extract, zhang2024dcpdd, maini2024dataset] flag the membership of $x$ in training by testing whether $S_{\theta}(x,x)$ is suspiciously low. We adopt the notation from Section 2.1: sequences $r,u\in V^{T}$, next-token marginals $p_{t}^{\,r},p_{t}^{\,u}$, their mixture $p_{t}^{\,d}=\alpha\,p_{t}^{\,r}+(1-\alpha)\,p_{t}^{\,u}$ with $\alpha=\frac{|r|}{|r|+|u|}$, and the mutual information $I(X_{\operatorname{MarI}};Z)$.
@@ 4 Experiments
GA [yao2024llmunlearning] ascent on unlearn set full information
GD [liu2022continual] ascent on unlearn set descent on retain set full information
KL-GA [hong2024dissect] ascent on unlearn set minimize $\mathrm{KL}\!\big(p_{\theta}^{\,r}\,\|\,p_{\theta_{0}}^{\,r}\big)$ full information
DPO [rafailov2023dpo] preference loss minimize $\mathrm{CE}$ or $\mathrm{KL}$ full information
@@ 4.1 Models, Datasets, and Splits
Protocol. Following maini2024tofutaskfictitiousunlearning, eldan2023approximate, we adopt a three-step evaluation: (i) fine-tune on $\mathcal{D}_{u}\!\cup\!\mathcal{D}_{r}$ to obtain a full finetune baseline; (ii) apply each unlearning method to remove the influence of $\mathcal{D}_{u}$ to obtain unlearn results, denoted by the name of corresponding unlearn methods; (iii) restart the fine-tune but only on $\mathcal{D}_{r}$ (never exposed to $\mathcal{D}_{u}$) to obtain an unlearn baseline. An ideal unlearning method should preserve retain/validation accuracy level similar to the unlearn baseline while lowering the unlearn accuracy to a similar level to the unlearn baseline. We report next-token accuracy on $\mathcal{D}_{r}$, $\mathcal{D}_{u}$, and a held-out validation set, and assess general capability with the Eleuther’s LM Evaluation Harness [eval-harness].
Datasets. We evaluate Forgetting–MarI on two mid-scale LMs, GPT-2 Large (774M) and Llama-3.2-1B, and two text domains with distinct genres and pretraining prevalence: (i) Harry Potter and the Prisoner of Azkaban [rowling2013harry], likely present in pretraining; and (ii) Careless People: A Cautionary Tale of Power, Greed, and Lost Idealism [wynn2025careless], published after Llama’s release and thus unlikely to appear in its pretraining.
Splits. For Harry Potter (HP), we designate 10% of sentences as $\mathcal{D}_{u}$ and the remaining 90% as $\mathcal{D}_{r}$ (cf. eldan2023approximate); validation uses excerpts from another book in the series (Harry Potter and the Sorcerer’s Stone [rowling1999harry]). For Careless People (CP), we consider: (i) a correlated split with contiguous 50/50 spans for $\mathcal{D}_{u}/\mathcal{D}_{r}$; and (ii) an uncorrelated split where $\mathcal{D}_{u}$ comprises 2025 Reddit stories (post-release) and $\mathcal{D}_{r}$ is 50% of the book. Reddit stories are used as validation in the first setting and the remaining 50% of the book is used as validation in the second setting.
@@ 4.2 Accuracy Trade-off & general capability
General capability (right panels). We compare the unlearned models against the finetuned baseline, the unlearn baseline, and the original GPT-2 (before finetuning) on ARC-E, PIQA, HellaSwag, WikiText, and MMLU [eval-harness].
@@ 4.4 Detector evaluation: Empirical verification of theoretical guarantees
Copyrighted-text detection methods fall into two families: (i) white-box detectors [carlini2021extract, zhang2024dcpdd, maini2024dataset, shi2023detecting], which use (tail or reference-model) perplexity to infer training membership; (ii) black-box detectors [karamolegkou2023leak, chang2023cloze, duarte2024decop, dong2024consistency, hu2024veilprobe], which rely on string-level similarity without logits. Because Forgetting-MarI (and most threat models) allow weight access, white-box tests are strictly harder to defeat; we thus focus on them. (See Appendix D for method details and additional results.)
We ran the current SOTA white-box detector [shi2023detecting] on: (i) the model finetuned on $\mathcal{D}_{r}\!\cup\!\mathcal{D}_{u}$, (ii) the gold-standard unlearn baseline trained only on $\mathcal{D}_{r}$, and (iii) model (i) after Forgetting-MarI. We report ROC–AUC: low values indicate the detector believes the model was trained with $\mathcal{D}_{u}$, high values indicate the opposite. As shown in Fig. 5, the ROC–AUC after Forgetting-MarI closely matches the unlearn baseline, indicating effective removal of $\mathcal{D}_{u}$’s influence, as predicted by theory.
@@ A.1 Details of LLM Unlearning Methods: Implicit Marginal Information Unlearning via conflicting forces
Recent surveys highlight four broad families of LLM-unlearning techniques, each making a different compromise between unlearn efficacy, the ability to remove information from a model, utility preservation, how well the model performs on the remaining data, and computational cost, the resources expended to perform the unlearning [liu2024rethinkingmachineunlearninglarge]. A heuristic commonality of the techniques is their implicit/indirect target of marginal unlearning: all the methods tend to detect and thereby remove only the marginal effect of adding an “unlearn set” ($\mathcal{D}_{u}$), the dataset that is meant to be forgotten, to a “retain set” ($\mathcal{D}_{r}$), the dataset that the model should remember, on the given model.
@@
These techniques train and perform weight updates on the whole model. Gradient ascent (or “loss reversal”) [yao2024llmunlearning] is the most straight-forward unlearning technique. It directly maximizes the cross-entropy on $\mathcal{D}_{u}$, effectively penalizing the model performance on the unlearn set. This type of unlearning was shown to lead to an overall decrease in model performance, so Gradient Difference [liu2024rethinkingmachineunlearninglarge] was developed to balance unlearning while maintaining general model performance. Gradient Difference maximizes the cross-entropy loss on $\mathcal{D}_{u}$ while continuing to minimize the loss on $\mathcal{D}_{r}$:
@@
Follow-up studies revealed that, even when balanced with gradient descent, this global ascent signal is too coarse: it suppresses the target examples but also degrades correlated yet legitimate content [hong2024dissect]. To overcome this challenge, variants have aimed to improve both sides of the problem. For utility preservation, past work has shown that distillation-style regularization with a Kullback–Leibler (KL) divergence penalty outperforms gradient descent on $\mathcal{D}_{r}$ in keeping the updated model close to the original without over-training on the retain set. For unlearning, alignment-style variants such as Direct Preference Optimization (DPO) and Negative Preference Optimization (NPO) replace the unlearn objective with more specific preference-based objectives, slowing catastrophic performance collapse [rafailov2023dpo, zhang2024npo]. However, such preference-supervised methods can be difficult to generalize to unlearn at a large scale.
@@
In an effort to perform unlearning more efficiently, this line of methods focuses on selectively altering only a subset of a model’s parameters rather than retraining the entire network. Such “model-surgery” methods perform rank-constrained updates at one or a few layers. Rank-One Model Editing (ROME) edits a single MLP weight with a closed-form rank-1 patch [meng2022rome]. In particular, it modifies only the weights causally responsible for one token sequence in the unlearn set:
@@
Here, $\Delta W:=\frac{\bigl(v_{\text{new}}-v_{\text{old}}\bigr)h^{\top}}{\|h\|_{2}^{2}}$, $l^{*}$ is the layer most influenced by the unlearn sample or prompt $x$, $h_{l^{*}}(x)$ is the activation and $W_{l^{*}}$ is the weight matrix of layer $l^{*}$, $v_{old}:=W_{l^{*}}h_{l^{*}}(x)$, and finally $v_{new}$ is the alternative answer we want to replace $v_{old}$ by. Mass Editing Memory in a Transformer (MEMIT) [meng2023memit] extends this idea to thousands of facts simultaneously and stacks many $(h^{i}_{l^{*}},v^{i}_{l^{*}})$ pairs. AlphaEdit furthers the idea by projecting edits into the null space of preserved knowledge, with the aim to improve robustness in sequential settings, ensuring minimal disruption to previously learned information. Detecting and Editing Privacy Neurons (DEPN) [wu2023depn] masks the gradients of neurons identified as contributing the most to the prediction of privacy-related content. In general, weight editing and partial tuning techniques are fast, but they are limited to short factual associations and struggle with stylistic or distributed knowledge.
@@
“I don’t know” [ishibashi2024sanitization] trains the model on question-answer pairs that map sensitive questions to a safe refusal (e.g. “I don’t know”), teaching the model to decline queries about the unlearn set. Entity anonymization [eldan2023approximate] replaces sensitive entities with anonymized placeholders and trains the model on the rewritten placeholders to scrub identifiable information from the model. Unlearning Large Language Models via Negative Response and Model Parameter Average (ULMR) [shi2024ulmr] constructs adversarial “negative” prompts, trains on the paired responses, and then averages the updated weights with the base model to dampen overshoot. Selective Knowledge‐negation Unlearning (SKU) first mines harmful or copyrighted contexts via red-teaming, then injects counterfactuals that negate them [liu2024sku]. Such approaches are easy to deploy but depend heavily on prompt engineering and high-quality counterexamples.
@@
These methods train something external to the model and then use that externally trained adapter to update the model itself. A common instantiation is the task-vector framework: let $p_{\theta_{0}}$ be the original model and $p_{\theta_{u}}$ the same model fine-tuned on the unlearn set $\mathcal{D}_{u}$. The element-wise difference $\Delta\theta:=\theta_{u}-\theta_{0}$ is treated as an encoding of the deleted knowledge and direct-subtraction methods [ilharco2023task] form the unlearned model as $p_{\theta_{0}-\Delta\theta}$. Orthogonality offers an alternative geometric control. $O^{3}$ [gao2024o3] trains one orthogonal LoRA adapter per removal request and learns a contrastive out-of-distribution (OOD) gate that activates the corresponding adapter at inference time. Orthogonality limits interference between requests, but the approach incurs two key costs: (1) the number of adapters (and hence memory) grows linearly with the number of unlearning requests, and (2) any mismatch between model behavior and the assumed linear/inner-product structure in weight space can undermine both unlearning guarantees and downstream utility.

### v3

@@ 1 Introduction
As machine learning models, particularly Large Language Models (LLMs), get trained on bigger datasets containing potentially sensitive or regulated information, and as LLMs are increasingly deployed in high-stakes domains, the need to selectively remove specific data influences from these models has become critical. This requirement is driven not only by privacy regulations such as the European Union’s General Data Protection Regulation (GDPR) and its “right to be forgotten,” but also by practical concerns including the removal of copyrighted content, personally identifiable information, or data determined to be harmful or biased [29, 13, 8, 4, 3]. Unlearning, or removing the influence of specific data post hoc, is an attractive tool for achieving this information removal, especially with the high costs of retraining a model from scratch.
@@ 1.1 Open Challenges in LLM Unlearning
Effective LLM unlearning must balance three objectives [23]. First, unlearn efficacy measures how well a model suppresses the influence of the data we want to unlearn, called the unlearn set $\mathcal{D}_{u}$. Second, utility preservation ensures the model’s ability to retain performance on general tasks and the data we are still authorized to use, called the retain set, $\mathcal{D}_{r}$ is not lost. Finally, computational cost encompasses the time, memory, and carbon used during unlearning. All unlearning techniques aim to optimize these three objectives, which inherently come with tradeoffs; what differs is where and how the model parameters are updated, directly affecting their ability to balance the three. A breakdown of existing techniques and their strengths and weaknesses is shown in Table 1, with their technical details and commonality in indirect marginal unlearning in Appendix A.1.
@@
Numbers map to BibTeX entries: 1[38], 2[23], 3[15], 4[31], 5[39], 7[27], 8[28], 9[7], 10[36], 11[18], 12[6], 13[33], 14[24], 15[17], 16[10].
@@
Robust unlearning & Utility Preservation: Existing LLM unlearning techniques via full-parameter fine-tuning typically treat the unlearn set $\mathcal{D}_{u}$ as fully toxic, forcing the model to forget every sequence in $\mathcal{D}_{u}$ regardless of their overlap with the retain set $\mathcal{D}_{r}$. Examples include loss-reversal [22], gradient-difference [38], KL-ascent [15], and preference-based DPO/NPO [31, 39]. Even local editors that aim to make precise edits (ROME, MEMIT) share this limitation [27, 28], erasing shared facts and stylistic cues, and raising perplexity on $\mathcal{D}_{r}$ and held-out tasks. Benchmarks (RWKU, MUSE, Eight-Method) consistently report sizable utility drops after unlearning [20, 35, 25].
Stable Continual Unlearning: As the legal landscape around data usage changes, a deployed LLM may receive hundreds or thousands of unlearn requests. Production-ready unlearning, therefore, needs to be able to repeatedly unlearn, retain utility, and keep computation and memory within a practical range. Exact methods like full retraining or shared SISA guarantee unlearning but their cost scales with both model size and request count [2, 12, 1]. Lighter updates like influence functions [14] or repeated ROME/MEMIT edits [27, 28] are cheap per removal yet accumulate inference costs and utility drift. Task-vector subtraction or adapter stacks save compute during unlearning but require storing external model adapters, also creating downstream inference costs [17, 10]. Thus, continually unlearning without runaway resources or utility loss remains unsolved.
Formal Guarantees at LLM-Scale: Certified unlearning is well established for linear/kernel models [14], high-dimensional classifiers [41], and general mathematical formulations of machine unlearning [37]. However, no existing method provides guarantees that scale to autoregressive transformers with billions of parameters (7B–70B+), such as GPT or Llama. As a result, practitioners lack reliable guarantees of the extent to which the unlearn set remains uninferable or undetectable after common downstream operations such as compression, distillation, or adversarial probing [23].
@@ 2.2 Marginal Information & Perplexity-Based Detectors
be the standard cross-entropy (per-token negative log-likelihood). State-of-the-art detectors [3, 40, 30] flag the membership of $x$ in training by testing whether $S_{\theta}(x,x)$ is suspiciously low. We adopt the notation from Section 2.1: sequences $r,u\in V^{T}$, next-token marginals $p_{t}^{\,r},p_{t}^{\,u}$, their mixture $p_{t}^{\,d}=\alpha\,p_{t}^{\,r}+(1-\alpha)\,p_{t}^{\,u}$ with $\alpha=\frac{|r|}{|r|+|u|}$, and the mutual information $I(X_{\operatorname{MarI}};Z)$.
@@ 4 Experiments
GA [38] ascent on unlearn set full information
GD [22] ascent on unlearn set descent on retain set full information
KL-GA [15] ascent on unlearn set minimize $\mathrm{KL}\!\big(p_{\theta}^{\,r}\,\|\,p_{\theta_{0}}^{\,r}\big)$ full information
DPO [31] preference loss minimize $\mathrm{CE}$ or $\mathrm{KL}$ full information
@@ 4.1 Models, Datasets, and Splits
Protocol. Following Maini et al. [26], Eldan and Russinovich [6], we adopt a three-step evaluation: (i) fine-tune on $\mathcal{D}_{u}\!\cup\!\mathcal{D}_{r}$ to obtain a full finetune baseline; (ii) apply each unlearning method to remove the influence of $\mathcal{D}_{u}$ to obtain unlearn results, denoted by the name of corresponding unlearn methods; (iii) restart the fine-tune but only on $\mathcal{D}_{r}$ (never exposed to $\mathcal{D}_{u}$) to obtain an unlearn baseline. An ideal unlearning method should preserve retain/validation accuracy level similar to the unlearn baseline while lowering the unlearn accuracy to a similar level to the unlearn baseline. We report next-token accuracy on $\mathcal{D}_{r}$, $\mathcal{D}_{u}$, and a held-out validation set, and assess general capability with the Eleuther’s LM Evaluation Harness [11].
Datasets. We evaluate Forgetting–MarI on two mid-scale LMs, GPT-2 Large (774M) and Llama-3.2-1B, and two text domains with distinct genres and pretraining prevalence: (i) Harry Potter and the Prisoner of Azkaban [19], likely present in pretraining; and (ii) Careless People: A Cautionary Tale of Power, Greed, and Lost Idealism [32], published after Llama’s release and thus unlikely to appear in its pretraining.
Splits. For Harry Potter (HP), we designate 10% of sentences as $\mathcal{D}_{u}$ and the remaining 90% as $\mathcal{D}_{r}$ (cf. 6); validation uses excerpts from another book in the series (Harry Potter and the Sorcerer’s Stone [rowling1999harry]). For Careless People (CP), we consider: (i) a correlated split with contiguous 50/50 spans for $\mathcal{D}_{u}/\mathcal{D}_{r}$; and (ii) an uncorrelated split where $\mathcal{D}_{u}$ comprises 2025 Reddit stories (post-release) and $\mathcal{D}_{r}$ is 50% of the book. Reddit stories are used as validation in the first setting and the remaining 50% of the book is used as validation in the second setting.
@@ 4.2 Accuracy Trade-off & general capability
General capability (right panels). We compare the unlearned models against the finetuned baseline, the unlearn baseline, and the original GPT-2 (before finetuning) on ARC-E, PIQA, HellaSwag, WikiText, and MMLU [11].
@@ 4.4 Detector evaluation: Empirical verification of theoretical guarantees
Copyrighted-text detection methods fall into two families: (i) white-box detectors [3, 40, 30, 34], which use (tail or reference-model) perplexity to infer training membership; (ii) black-box detectors [21, 9, 5, 16], which rely on string-level similarity without logits. Because Forgetting-MarI (and most threat models) allow weight access, white-box tests are strictly harder to defeat; we thus focus on them. (See Appendix D for method details and additional results.)
We ran the current SOTA white-box detector [34] on: (i) the model finetuned on $\mathcal{D}_{r}\!\cup\!\mathcal{D}_{u}$, (ii) the gold-standard unlearn baseline trained only on $\mathcal{D}_{r}$, and (iii) model (i) after Forgetting-MarI. We report ROC–AUC: low values indicate the detector believes the model was trained with $\mathcal{D}_{u}$, high values indicate the opposite. As shown in Fig. 5, the ROC–AUC after Forgetting-MarI closely matches the unlearn baseline, indicating effective removal of $\mathcal{D}_{u}$’s influence, as predicted by theory.
@@ A.1 Details of LLM Unlearning Methods: Implicit Marginal Information Unlearning via conflicting forces
Recent surveys highlight four broad families of LLM-unlearning techniques, each making a different compromise between unlearn efficacy, the ability to remove information from a model, utility preservation, how well the model performs on the remaining data, and computational cost, the resources expended to perform the unlearning [23]. A heuristic commonality of the techniques is their implicit/indirect target of marginal unlearning: all the methods tend to detect and thereby remove only the marginal effect of adding an “unlearn set” ($\mathcal{D}_{u}$), the dataset that is meant to be forgotten, to a “retain set” ($\mathcal{D}_{r}$), the dataset that the model should remember, on the given model.
@@
These techniques train and perform weight updates on the whole model. Gradient ascent (or “loss reversal”) [38] is the most straight-forward unlearning technique. It directly maximizes the cross-entropy on $\mathcal{D}_{u}$, effectively penalizing the model performance on the unlearn set. This type of unlearning was shown to lead to an overall decrease in model performance, so Gradient Difference [23] was developed to balance unlearning while maintaining general model performance. Gradient Difference maximizes the cross-entropy loss on $\mathcal{D}_{u}$ while continuing to minimize the loss on $\mathcal{D}_{r}$:
@@
Follow-up studies revealed that, even when balanced with gradient descent, this global ascent signal is too coarse: it suppresses the target examples but also degrades correlated yet legitimate content [15]. To overcome this challenge, variants have aimed to improve both sides of the problem. For utility preservation, past work has shown that distillation-style regularization with a Kullback–Leibler (KL) divergence penalty outperforms gradient descent on $\mathcal{D}_{r}$ in keeping the updated model close to the original without over-training on the retain set. For unlearning, alignment-style variants such as Direct Preference Optimization (DPO) and Negative Preference Optimization (NPO) replace the unlearn objective with more specific preference-based objectives, slowing catastrophic performance collapse [31, 39]. However, such preference-supervised methods can be difficult to generalize to unlearn at a large scale.
@@
In an effort to perform unlearning more efficiently, this line of methods focuses on selectively altering only a subset of a model’s parameters rather than retraining the entire network. Such “model-surgery” methods perform rank-constrained updates at one or a few layers. Rank-One Model Editing (ROME) edits a single MLP weight with a closed-form rank-1 patch [27]. In particular, it modifies only the weights causally responsible for one token sequence in the unlearn set:
@@
Here, $\Delta W:=\frac{\bigl(v_{\text{new}}-v_{\text{old}}\bigr)h^{\top}}{\|h\|_{2}^{2}}$, $l^{*}$ is the layer most influenced by the unlearn sample or prompt $x$, $h_{l^{*}}(x)$ is the activation and $W_{l^{*}}$ is the weight matrix of layer $l^{*}$, $v_{old}:=W_{l^{*}}h_{l^{*}}(x)$, and finally $v_{new}$ is the alternative answer we want to replace $v_{old}$ by. Mass Editing Memory in a Transformer (MEMIT) [28] extends this idea to thousands of facts simultaneously and stacks many $(h^{i}_{l^{*}},v^{i}_{l^{*}})$ pairs. AlphaEdit furthers the idea by projecting edits into the null space of preserved knowledge, with the aim to improve robustness in sequential settings, ensuring minimal disruption to previously learned information. Detecting and Editing Privacy Neurons (DEPN) [36] masks the gradients of neurons identified as contributing the most to the prediction of privacy-related content. In general, weight editing and partial tuning techniques are fast, but they are limited to short factual associations and struggle with stylistic or distributed knowledge.
@@
“I don’t know” [18] trains the model on question-answer pairs that map sensitive questions to a safe refusal (e.g. “I don’t know”), teaching the model to decline queries about the unlearn set. Entity anonymization [6] replaces sensitive entities with anonymized placeholders and trains the model on the rewritten placeholders to scrub identifiable information from the model. Unlearning Large Language Models via Negative Response and Model Parameter Average (ULMR) [33] constructs adversarial “negative” prompts, trains on the paired responses, and then averages the updated weights with the base model to dampen overshoot. Selective Knowledge‐negation Unlearning (SKU) first mines harmful or copyrighted contexts via red-teaming, then injects counterfactuals that negate them [24]. Such approaches are easy to deploy but depend heavily on prompt engineering and high-quality counterexamples.
@@
These methods train something external to the model and then use that externally trained adapter to update the model itself. A common instantiation is the task-vector framework: let $p_{\theta_{0}}$ be the original model and $p_{\theta_{u}}$ the same model fine-tuned on the unlearn set $\mathcal{D}_{u}$. The element-wise difference $\Delta\theta:=\theta_{u}-\theta_{0}$ is treated as an encoding of the deleted knowledge and direct-subtraction methods [17] form the unlearned model as $p_{\theta_{0}-\Delta\theta}$. Orthogonality offers an alternative geometric control. $O^{3}$ [10] trains one orthogonal LoRA adapter per removal request and learns a contrastive out-of-distribution (OOD) gate that activates the corresponding adapter at inference time. Orthogonality limits interference between requests, but the approach incurs two key costs: (1) the number of adapters (and hence memory) grows linearly with the number of unlearning requests, and (2) any mismatch between model behavior and the assumed linear/inner-product structure in weight space can undermine both unlearning guarantees and downstream utility.